OpenStudy (anonymous):

Given the linear equation y=3/4x-2, find the y-coordinates of the points (-8, ), (-4, ), and (4, ). Please show all of your work

OpenStudy (anonymous):

put the x value you will get corresponding y.

OpenStudy (anonymous):

if (8, ) is x value then y value will be (put x=8 in the given equation) \[\Large y=\frac{3}{4}(8)-22=\frac{24}{4}-2=6-2=4\] so y value is 4can you find it now ?

OpenStudy (anonymous):

@amber517 did u find it?

OpenStudy (anonymous):

yes thanks